# TokLIP: Marry Visual Tokens to CLIP for Multimodal Comprehension and Generation

### v1

@@ 4.5 Ablation Studies
We explore several training strategies for our TokLIP models. Specifically, we evaluate three key aspects: (1) whether to initialize the vision encoder with pre-trained weights, (2) the method for projecting VQ codes into the CLIP feature space, and (3) the design of the learning objectives. To analyze these strategies, we implement various ablation models based on OpenCLIP ViT-B/16 [9] and pre-train on CC3M dataset [5], listing the results in Table 5.
@@
Bidirectional TokLIP.
As introduced in Section 3.1, we apply a causal attention module in TokLIP for unified understanding and generation. Our causal attention-based TokLIP demonstrates superior semantic understanding compared to the VILA-U. In this section, we implement a bidirectional attention version using OpenCLIP-ViT-B/16 [9] to investigate whether a bidirectional architecture could further enhance comprehension tasks. We pre-train both the causal and bidirectional models on the CC3M and CC12M datasets with the same training setups. Results in Table 6 show that bidirectional-based models surpass causal-based models on classification and retrieval tasks. This suggests that a bidirectional TokLIP can further improve comprehension capacity for MLLMs, reinforcing the idea that learning semantics from discrete tokens is more effective than discretizing continuous high-level features.
Table 6: Ablations on bidirectional and causal attention in TokLIP.
Vision Encoder Attention Dataset IN Top1$\uparrow$ COCO TR@1$\uparrow$ COCO IR@1$\uparrow$
ViT-B/16 Bidirectional CC3M 56.32 44.66 32.66
ViT-B/16 Causal CC3M 52.46 41.30 30.13
ViT-B/16 Bidirectional CC12M 58.99 47.32 33.78
ViT-B/16 Causal CC12M 55.02 42.80 31.06
@@ B.1 Further Discussion
Vision Encoder Attention Distillation IN Top1$\uparrow$ COCO TR@1$\uparrow$ COCO IR@1$\uparrow$

### v2

@@ 4.5 Ablation Studies
We explore several training strategies for our TokLIP models. Specifically, we evaluate three key aspects: (1) whether to initialize the vision encoder with pre-trained weights, (2) the method for projecting VQ codes into the CLIP feature space, and (3) the design of the learning objectives. To analyze these strategies, we implement various ablation models based on OpenCLIP ViT-B/16 (Cherti et al., 2023) and pre-train on CC3M dataset (Changpinyo et al., 2021), listing the results in Table 5.
@@
4.6 Fine-grained Evaluation
To further assess the generalization capacity of the TokLIP tokenizer, we conduct detailed evaluations on fine-grained zero-shot classification tasks. We compare TokLIP with QLIP (Zhao et al., 2025) across nine widely used benchmarks, including Caltech101 (Fei-Fei et al., 2004), Flowers102 (Nilsback and Zisserman, 2008), Oxford Pets (Parkhi et al., 2012), Stanford Cars (Krause et al., 2013), FGVC Aircraft (Maji et al., 2013), SUN397 (Xiao et al., 2010), and Food101 (Bossard et al., 2014). As shown in Table 7, TokLIP consistently outperforms QLIP on all datasets, achieving a higher average accuracy (85.03 vs. 78.88), highlighting its strong generalization. These results further validate the effectiveness of our semanticization pipeline for VQ tokens.
Table 7: TokLIP performance on fine-grained classification benchmarks.
Model Res. Pets Flowers Food101 SUN397 Cars CalTech101 Aircraft Avg.
QLIP 392 91.03 78.63 89.66 72.29 88.25 97.32 34.95 78.88
TokLIP 384 93.64 88.58 89.98 74.77 93.31 97.65 57.25 85.03
Considering fine-grained capacity in MLLMs, TokLIP demonstrates a strong ability to understand scientific diagrams, as evidenced by its superior performance on the AI2D dataset. However, VQ tokens inherently suffer from information loss due to quantization, which is a well-known limitation for discrete MLLMs (Xie et al., 2024a; Team, 2024; Wang et al., 2024b). We freeze the VQ encoder to ensure extensibility to stronger VQ tokenizers, but this design may cause a certain performance drop when dealing with text-rich figures, where fine-grained text information is small. This limitation mainly originates from the VQ tokenizer itself (Shi et al., 2024a; Agarwal et al., 2025). Several works (Li et al., 2024b; Xie et al., 2025; Wang et al., 2024b, 2025b) have explored ways to mitigate this issue, such as training with high-resolution images, utilizing text-rich datasets, or unlocking the VQ tokenizers. Reinforcement learning has also proven to be an effective method for enhancing the capabilities of unified MLLMs in both fine-grained generation and understanding tasks (Geng et al., 2025; Xiao et al., 2025; Jiang et al., 2025). TokLIP can be further enhanced by adopting these strategies, and we plan to explore them in the future work.
4.7 Bidirectional vs. Causal Attention
As introduced in Section 3.1, we apply a causal attention module in TokLIP for unified understanding and generation. Our causal attention-based TokLIP demonstrates superior semantic understanding compared to the VILA-U. In this section, we present a detailed analysis of bidirectional and causal attention mechanisms, examining their respective impacts on comprehension and generation tasks.
Bidirectional attention enhances comprehension.
We first implement a bidirectional attention version using OpenCLIP-ViT-B/16 (Cherti et al., 2023) to investigate whether a bidirectional architecture could further enhance comprehension tasks. We pre-train both the causal and bidirectional models on the CC3M and CC12M datasets with the same training setups. Results in Table 8 show that bidirectional-based models surpass causal-based models on classification and retrieval tasks. This suggests that a bidirectional TokLIP can further improve comprehension capacity for MLLMs, reinforcing the idea that learning semantics from discrete tokens is more effective than discretizing continuous features.
Table 8: Ablations on bidirectional and causal attention in TokLIP.
[…]
ViT-B/16 Bidirectional CC3M 56.32 44.66 32.66
ViT-B/16 Causal CC3M 52.46 41.30 30.13
ViT-B/16 Bidirectional CC12M 58.99 47.32 33.78
ViT-B/16 Causal CC12M 55.02 42.80 31.06
Figure 5: Illustration of non-causal info for bidirectional TokLIP and quantitative generation results.
Causal attention underpins generation.
Causal TokLIP is necessary for generation tasks: 1) TokLIP with bidirectional attention introduces non-causal information flow as shown in Figure 5, which breaks the autoregressive nature for next-token prediction. 2) We conduct ablation of causal vs. bidirectional TokLIP. Both models are trained on CC3M using the LlamaGen-B framework. Bidirectional TokLIP (FID 214.36 on ImageNet) performs significantly worse than causal one (15.05) and baseline VQGAN (17.31). Performance degradation lies in the gap between non-causal training and autoregressive inference, which confirms that causal attention is essential.
Based on the detailed analysis and experimental results on both comprehension and generation tasks, we conclude that causal attention is the preferred choice for TokLIP. This decision is motivated by its ability to preserve autoregressive generation capability, which is essential for maintaining a unified architecture.
@@ B.1 Further Discussion
[…]
@@
B.2 Effects of TokLIP with small LLMs.
To further evaluate the effectiveness of TokLIP, we train TokLIP-XL with a smaller LLM, Qwen-2.5-1.5B, using the LLaVA-v1.5 dataset. We compare our model with MAR (Wu et al., 2025), which uses the same LLM but is trained on a significantly larger dataset (29.8M vs. 1.2M). As shown in Table B3, TokLIP achieves competitive results compared to MAR, even with much fewer data sources. This showcases the effectiveness of TokLIP even when paired with smaller-scale language models.
Table B3: Comprehension results with Qwen2.5-1.5B-Instruct.
Model Data POPE MME-P MMB SEED MMMU
MAR-H 25M pt+Onevison sft 87.6 1155.0 65.5 67.1 38.9
TokLIP LLaVA-v1.5 84.4 1222.0 64.5 64.7 39.4